# Where did heavy binaries go? Gravitational-wave populations using Delaunay triangulation with optimized complexity

### v1

@@ 4 Conclusions
We presented a non-parametric reconstruction of the mass-redshift correlation in merging BBH using Delaunay triangulation and barycentric interpolation. Our analysis of GWTC-4.0 data reveals that the high-mass peak at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$ appears predominantly from $z\gtrsim 0.7$ and is absent at $z=0.2$.
So, where would the heavy binaries have gone? While a detailed comparison with astrophysical predictions is necessary, interpreting our results will require models in which BBH sub-populations contribute to the observed merger rate at different epochs, possibly shaped by the combined influence of stellar evolution, metallicity-dependent mass loss, and the relative importance of dynamical formation channels.
Moreover, our findings imply that analyses measuring cosmological parameters from the redshift dependence of the detector-frame mass spectrum (Chernoff & Finn, 1993; Markovic, 1993; Farr et al., 2019; Mastrogiovanni et al., 2021; Abac et al., 2025a) should carefully account for the evolution of the prominence of the high-mass peak with $z$.
Our findings suggest a correlation between BBH mass and merger redshift in GWTC-4.0 data. At the same time, we caution against overinterpreting (the medians of) non-parametric reconstructions, which often, including here, come with large statistical errors.
The key strength of using Delaunay triangulation is that the interpolation skeleton, including the number and location of the nodes, is inferred directly from the data. In practice, this optimized complexity reduces the number of parameters by one to two orders of magnitude compared to the pixelized approach by Heinzel et al. (2025), and attain a good recovery of arbitrary functional forms at a lower tuning cost compared to the binned Gaussian-process approach by Ray et al. (2023), both of which rely on a fixed grid. For this reason, we expect our flexible framework to be significantly more efficient at probing correlations in dimensions $\geq 3$ (as we will illustrate in an upcoming publication; Toubiana et al. 2025a), making it ideally suited for the forthcoming big-data era of GW astronomy.
We thank April Qiu Cheng, Raffi Enficiaud, Jack Heinzel, and Matthew Mould for discussions. R.T., A.T., T.B. and D.G. ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, Italian-French University (UIF/UFI) Grant No. 2025-C3-386, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. A.T. and D.G. are supported by MUR Young Researchers Grant No. SOE2024-0000125. D.G. is supported by MSCA Fellowship No. 101149270–ProtoBH. Computational work was performed at CINECA with allocations through INFN and the University of Milano-Bicocca, and at NVIDIA with allocations through the Academic Grant program. This research has made use of data or software obtained from the Gravitational Wave Open Science Center.
@@ Appendix A Delaunay triangulation for inference
where $|S|$ denotes the generalized volume of $S$ (i.e. the area if $D=2$) and $w_{1,\dots,D+1}$ are the weights of its vertices.
@@
Appendix B Formalism for hierarchical inference
@@
$p(\{d_{j}\}|\Lambda)\propto e^{-N_{\mathrm{det}}(\Lambda)}\prod_{j=1}^{N_{\mathrm{obs}}}\left\langle\frac{\mathrm{d}_{\theta}N(\theta|\Lambda)}{\pi_{\mathrm{PE}}(\theta)}\right\rangle_{\theta\sim p(\theta|d_{j})}\,$, (B1)
@@
$N_{\rm det}(\Lambda)=\left\langle\frac{\mathrm{d}_{\theta}N(\theta|\Lambda)}{\pi_{\mathrm{inj}}(\theta)}p(\mathrm{det}|\theta)\right\rangle_{\theta\sim\pi_{\mathrm{inj}}(\theta)}$ (B2)
@@
$p({\rm det}|\theta)=\int_{d>\text{threshold}}p(d|\theta)\,{\rm d}d,$ (B3)
@@
Equations (B1) and (B2) are computed via Monte Carlo integration, here indicated with $\langle\cdot\rangle$. Selection effects are estimated using the injection campaign presented by Essick et al. (2025). We set the population likelihood to zero for hyperparameters that yield a variance in the Monte Carlo estimators greater than one (Talbot & Golomb, 2023; Abac et al., 2025b).
@@
Appendix C Population inference for $(m_{1},z)$
To facilitate comparison with other studies, Fig. 4 shows the results of our method applied to the joint $(m_{1},z)$ distribution instead of $(M_{\mathrm{tot}},z)$. The significance of the enhanced rate at $\sim 35\,\mathrm{M}_{\odot}$ at $z=1$ versus $z=0.2$ is slightly lower than for $(M_{\mathrm{tot}},z)$ ($95\%$ vs. $98.5\%$, right panel of Fig. 4), likely due to larger measurement uncertainties in $m_{1}$, but the qualitative disappearance of the high-mass peak at low redshift persists.

### v2

@@ 4 Conclusions
We presented a non-parametric reconstruction of the mass-redshift correlation in merging BBH using Delaunay triangulation and barycentric interpolation. Our analysis of GWTC-4.0 data reveals that the high-mass feature at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$ appears predominantly from $z\gtrsim 0.7$ and is absent at $z=0.2$. At the same time, we caution against overinterpreting (the medians of) non-parametric reconstructions, which often, including here, come with large statistical errors.
So, where would the heavy binaries have gone? Taken together, our results favor a mixed population at $M_{\mathrm{tot}}\sim 70\,\mathrm{M}_{\odot}$ with at least two components: a short–time-delay channel stemming from dense environments that dominates at high redshift, and a longer–time-delay component consistent with isolated (and ejected) binaries that becomes dominant toward low redshift.
Moreover, our findings imply that analyses measuring cosmological parameters from the redshift dependence of the detector-frame mass spectrum (Chernoff & Finn, 1993; Markovic, 1993; Farr et al., 2019; Mastrogiovanni et al., 2021; Abac et al., 2025b) should carefully account for the evolution of the prominence of the high-mass peak with $z$.
The key strength of using Delaunay triangulation is that the interpolation skeleton, including the number and location of the nodes, is inferred directly from the data. In practice, this optimized complexity reduces the number of parameters by one to two orders of magnitude compared to the pixelized approach by Heinzel et al. (2025b), and attain a good recovery of arbitrary functional forms at a lower tuning cost compared to the binned Gaussian-process approach by Ray et al. (2023), both of which rely on a fixed grid. For this reason, we expect our flexible framework to be significantly more efficient at probing correlations in dimensions $\geq 3$, making it ideally suited for the forthcoming big-data era of GW astronomy.
We thank April Qiu Cheng, Raffi Enficiaud, Jack Heinzel, and Matthew Mould for discussions. R.T., A.T., T.B. and D.G. are supported by ERC Starting Grant No. 945155–GWmining, Cariplo Foundation Grant No. 2021-0555, MUR PRIN Grant No. 2022-Z9X4XS, Italian-French University (UIF/UFI) Grant No. 2025-C3-386, MUR Grant “Progetto Dipartimenti di Eccellenza 2023-2027” (BiCoQ), and the ICSC National Research Centre funded by NextGenerationEU. A.T. and D.G. are supported by MUR Young Researchers Grant No. SOE2024-0000125. D.G. is supported by MSCA Fellowship No. 101149270–ProtoBH. Computational work was performed at CINECA with allocations through INFN and the University of Milano-Bicocca, and at NVIDIA with allocations through the Academic Grant program. This research has made use of data or software obtained from the Gravitational Wave Open Science Center.
@@ Appendix A Delaunay triangulation for inference
where $|S|$ denotes the generalized volume of $S$ (i.e. the area if $D=2$) and $w_{1},\dots,w_{D+1}$ are the weights of its vertices.
@@
Appendix B Mass-ratio distribution
The fiducial model by Abac et al. (2025c) describes the joint $(m_{1},q)$ distribution as a sum of components, each independent in $m_{1}$ and $z$. This formulation makes the overall distribution non-separable in $m_{1}$ and $q$, and this cannot be straightforwardly implemented in our framework. Guided by Abac et al. (2025c)’s findings (see their Fig. 5), we describe the mass-ratio distribution using a broken power law:
$p(q|\Lambda_{q})\propto\begin{cases}0&q<q_{\mathrm{min}}\,,\\ (q/q_{\mathrm{cut}})^{\alpha_{1}}&q_{\mathrm{min}}<q<q_{\mathrm{cut}}\,,\\ (q/q_{\mathrm{cut}})^{\alpha_{2}}&q_{\mathrm{cut}}\leq q\,.\end{cases}\,.$ (B1)
Prior distributions for $\Lambda_{q}=(q_{\mathrm{cut}},\alpha_{1},\alpha_{2})$ are uniform across $(0.5,1.0)$, $(0,10)$, and (10,10)(-10,10) respectively. The lower end $q_{\mathrm{min}}=3\,\mathrm{M}_{\odot}/m_{1}$ is chosen so that $m_{2}>3\,\mathrm{M}_{\odot}$. The population predictive distribution of $q$ and the hyperposterior distribution of $(\alpha_{1},\alpha_{2})$ are shown in Fig. 5. We observe an enhanced number of mergers at $0.6\lesssim q\lesssim 0.8$ compared to the single power-law model.
For completeness, we also show the results for a single power law with, which corresponds to $q_{\mathrm{cut}}\rightarrow 1$ and/or $\alpha_{2}\to\alpha_{1}$. The conclusions of our work regarding the behavior of the merger rate across $(M_{\mathrm{tot}},z)$ are overall unchanged. The main quantitative difference is that the credibility for an enhanced rate at $\sim 70\,\mathrm{M}_{\odot}$ is lower ($\sim 94\%$ vs. $\sim 98\%$).
Figure 5: Left: Posterior predictive distribution of the mass-ratio distribution. The red (grey) solid line represent the median, while the shade (dotted) envelope contains the 90% credible interval assuming a broken (single) power-law distribution. Right: Posterior distribution of the power-law indices. The gray area represents the posterior probability on $(\alpha_{1},\alpha_{2})$ for the broken power-law model, while the dotted histogram represents the posterior probability on $\alpha_{1}$ for the single power-law model.
Appendix C Formalism for hierarchical inference
@@
$p(\{d_{j}\}|\Lambda)\propto e^{-N_{\mathrm{det}}(\Lambda)}\prod_{j=1}^{N_{\mathrm{obs}}}\left\langle\frac{\mathrm{d}_{\theta}N(\theta|\Lambda)}{\pi_{\mathrm{PE}}(\theta)}\right\rangle_{\theta\sim p(\theta|d_{j})}\,$, (C1)
@@
$N_{\rm det}(\Lambda)=\left\langle\frac{\mathrm{d}_{\theta}N(\theta|\Lambda)}{\pi_{\mathrm{inj}}(\theta)}p(\mathrm{det}|\theta)\right\rangle_{\theta\sim\pi_{\mathrm{inj}}(\theta)}$ (C2)
@@
$p({\rm det}|\theta)=\int_{d>\text{threshold}}p(d|\theta)\,{\rm d}d,$ (C3)
@@
Equations (C1) and (C2) are computed via Monte Carlo integration, here indicated with $\langle\cdot\rangle$. Selection effects are estimated using the injection campaign presented by Essick et al. (2025). We set the population likelihood to zero for hyperparameters that yield a variance in the Monte Carlo estimators greater than one (Talbot & Golomb, 2023; Abac et al., 2025c; Mancarella & Gerosa, 2025). We verify the robustness of this choice by re-running our inference using thresholds of 0.8 and 1.2. The results are qualitatively unchanged; the credibility for an enhanced merger rate at $\sim 70\,\mathrm{M}_{\odot}$ is $\sim 95\%$ instead of $\sim 98\%$ for the 1.2 threshold.
As discussed in the main text, the prior on the modeled part of the population parameters $(\Lambda_{q},\Lambda_{\chi},\Lambda_{\vartheta})$ are taken equal to those of Abac et al. (2025c), while the prior on the triangulation parameters is uniform across the area of interest for the triangulation vertices and along (20,15)(-20,15) for the triangulation weights. The implied prior range on $\mathrm{d}_{\theta}N$ (or equivalently $\mathrm{d}_{\theta}\mathcal{R}$) is compatible with (if not more conservative than) those used in similar non-parametric approaches (Heinzel et al., 2025b, a; Alvarez-Lopez et al., 2025). We also verified the robustness of our results against prior choices by extending the weight prior down to (25,15)(-25,15), which lowers the induced prior on $\mathrm{d}_{\theta}N$ by two to three orders of magnitude. The results for $z=0.2$ remain unchanged. For $z=1$, the lower end of the posterior distribution on $\mathrm{d}_{\theta}\mathcal{R}$ lowers so that the credibility for $\mathrm{d}_{\theta}\mathcal{R}(z=1)>\mathrm{d}_{\theta}\mathcal{R}(z=0.2)$ is slightly lower ($\sim 94\%$ instead of $\sim 98\%$). We stress that the key result of this work, namely the disappearance of the peak at low redshifts, is not affected by prior choices. The lack of an obvious choice for the prior range of $w_{i}$ suggests non-parametric population models are better suited to conduct inference in the observed space, where the domain of inference is naturally limited to data-dominated regions (Toubiana et al., 2025).
We also compute the Kullback-Leibler divergence $D_{\mathrm{KL}}$ (Kullback & Leibler, 1951) to compare the posterior volumetric merger rate to its prior distribution at any given $(M_{\mathrm{tot}},z)$. The prior distribution on $\log_{10}\mathrm{d}_{\theta}\mathcal{R}(M_{\mathrm{tot}},z)$, denoted by $p(\log_{10}\mathrm{d}_{\theta}\mathcal{R})$ for simplicity, is constructed by sampling the prior distribution of $\Lambda_{\Delta}$ (i.e. number of vertices, positions, and weights) and computing the corresponding $\log_{10}\mathrm{d}_{\theta}\mathcal{R}(M_{\mathrm{tot}},z)$. The posterior distribution, denoted $p(\log_{10}\mathrm{d}_{\theta}\mathcal{R}|\{d_{j}\})$, is computed analogously by sampling from $p(\Lambda_{\Delta}|\{d_{j}\})$. To approximate these probability densities, we generate $10^{4}$ samples in every $(M_{\mathrm{tot}},z)$ of interest and fit a Gaussian Kernel Density Estimator. The Kullback-Leibler divergence (in nat) for any given $(M_{\mathrm{tot}},z)$ is then computed as
$D_{\mathrm{KL}}=\left\langle\ln\frac{p(\log_{10}\mathrm{d}_{\theta}\mathcal{R}|\{d_{j}\})}{p(\log_{10}\mathrm{d}_{\theta}\mathcal{R})}\right\rangle_{\log_{10}\mathrm{d}_{\theta}\mathcal{R}\sim p(\log_{10}\mathrm{d}_{\theta}\mathcal{R}|\{d_{j}\})}\,.$ (C4)
Appendix D Population inference for $(m_{1},z)$
[…]